Apply the method of variations of parameters to solve the following differential equations:
1) x^2 y" + xy' - y = x^2 e^x
2) y" + a^2 y = cosec ax
3) solve the equation d^2 y/dx^2 - cotx dy/dx - sin^2 xy = cosx - cos^3 x by changing the independent variable
1
Expert's answer
2020-03-12T12:33:17-0400

1) "x^2y''+xy'-y=x^2e^x"

We will replace

"x=e^t, t=ln|x|\\
y'=\frac{dy}{dx}=\frac{y'_t}{x'_t}=\frac{y'_t}{e^t}=e^{-t}y'_t\\
y''=\frac{dy'}{dx}=\frac{(e^{-t}\cdot y'_{t})'}{e^t}=\frac{y''_{t}e^{-t}-e^{-t}y'_t}{e^t}=e^{-2t}(y''_t-y'_t)"

then

"e^{2t}e^{-2t}(y''-y')+e^te^{-t}y'-y=e^{2t}\cdot e^{e^t}\\
y''-y'+y'-y=e^{2t}\cdot e^{e^t}\\
y''-y=e^{2t}\cdot e^{e^t}"

a)

"y''-y=0\\
\lambda^2-1=0\\
\lambda_1=1, \lambda_2=-1\\
y_1=e^t, y_2=e^{-t}\\
y_0=c_1e^t+c_2e^{-t}"

b)

"y=c_1(t)e^t+c_2(t)e^{-t}\\"


"\left\{\begin{matrix}
 c'_1e^t+c'_2e^{-t}=0 \\
 c'_1e^t-c'_2e^{-t}=e^{2t}e^{e^t}
\end{matrix}\right.\\
2c'_1e^t=e^{2t}e^{e^t}\\
c'_1=\frac{1}{2}e^{t}e^{e^t}\\
c_1=\frac{1}{2}\int e^{t}e^{e^t}dt=\frac{1}{2}\int e^{e^t}d(e^t)=\\
=\frac{1}{2}e^{e^t}+k_1"


"2c'_2e^{-t}=-e^{2t}e^{e^t}\\
c'_2=-\frac{1}{2}e^{3t}e^{e^t}\\
c_2=-\frac{1}{2}\int e^{3t}e^{e^t}dt=-\frac{1}{2}\int e^{2t}e^{e^t}d(e^t)"

let "u=e^t" then

"\int e^{2t}e^{e^t}de^t=\int u^2e^udu=u^2e^u-\int2ue^udu=\\
=u^2e^u-2(ue^u-\int e^udu)=u^2e^u-2ue^u+2e^u\\
c_2=-\frac{1}{2}(u^2e^u-2ue^u+2e^u)+k_2=\\
=-\frac{1}{2}(e^{2t}e^{e^t}-2e^te^{e^t}+2e^{e^t})+k_2"


Then

"y(t)=(\frac{1}{2}e^{e^t}+k_1)e^t+\\
+(-\frac{1}{2}(e^{2t}e^{e^t}-2e^te^{e^t}+2e^{e^t})+k_2)e^{-t}\\
y(x)=(\frac{1}{2}e^x+k_1)x+\\
+(-\frac{1}{2}(x^2e^x-2xe^x+2e^x)+k_2)\frac{1}{x}\\"


2)

"y''+a^2y=cosecax=\frac{1}{sinax}"

a)

"y''+a^2y=0\\
\lambda^2+a^2=0\\
\lambda_1=ai, \lambda_2=-ai\\
y_1=cosax, y_2=sinax\\
y_0=c_1cosax+c_2sinax"


b)

"y=c_{1}(x)cosax+c_2(x)sinax\\
\left\{\begin{matrix}
 c'_1cosax+c'_2 sinax=0 \\
 -a c'_1sinax+ac'_2 cosax=\frac{1}{sinax} 
\end{matrix}\right."


"\Delta=\begin{vmatrix}
 cos ax & sinax\\
 -asinax & acosax
\end{vmatrix}=a\\
\Delta_1=\begin{vmatrix}
 0 & sinax\\
 \frac{1}{sinax} & acosax
\end{vmatrix}=-1\\
\Delta_2=\begin{vmatrix}
 cos ax & 0\\
 -asinax & \frac{1}{sinax}
\end{vmatrix}=\frac{cosax}{sinax}\\
c'_{1}=\frac{\Delta_1}{\Delta}=\frac{-1}{a}\\
c_{1}=\int\frac{-1}{a}dx=-\frac{x}{a}+k_1\\
c'_2=\frac{\Delta_2}{\Delta}=\frac{1}{a}\cdot\frac{cosax}{sinax}\\
c_2=\int\frac{1}{a}\cdot\frac{cosax}{sinax}dx=\frac{1}{a^2}ln|sinax|+k_2"

"y=(-\frac{x}{a}+k_1)cosax+(\frac{1}{a^2}ln|sinax|+k_2)sinax"


3) "y''-\cot x y'- \sin^2x y=\cos x-\cos^3x"

We will replace

"\cos x=t, x=\arccos t, \sin x =\sqrt{1-t^2}\\
y'=\frac{dy}{dx}=\frac{y'_t}{x'_t}=\frac{y'_t}{-\frac{1}{\sqrt{1-t^2}}}=-\sqrt{1-t^2} y'_t\\
y''=\frac{dy'}{dx}=\frac{-y''_t\sqrt{1-t^2}-y'_t\frac{1}{2\sqrt{1-t^2}}(-2t)}{-\frac{1}{\sqrt{1-t^2}}}=\\
=y''_t(1-t^2)-2ty'_t"

input in equation

"y''(1-t^2)-ty'-\frac{t}{\sqrt{1-t^2}}(-\sqrt{1-t^2})y'-\\
-(1-t^2)y=t-t^3\\
y''(1-t^2)-(1-t^2)y=t(1-t^2)\\
y''-y=t,\\
y=y_0+y_1\\
a) y''-y=0\\
\lambda^2-1=0, \lambda_1=1, \lambda_2=-1\\
y_1=e^t, y_2=e^{-t}\\
y_0=c_1e^t+c_2e^{-t}\\
b) y_1=at+b"

input in equation

"-at-b=t\implies a=-1, b=0,\\
y_1=-t\\
y=c_1e^t+c_2e^{-t}-t\\
y=c_1e^{\cos x}+c_2e^{-\cos x}-\cos x"